Now the student moves the box up a ramp
(with the same coefficient of friction) inclined
at 10.1
◦ with the horizontal.
b) If the box starts from rest at the bottom
of the ramp and is pulled at an angle of 26.7

with respect to the incline and with the same
173 N force, what is the acceleration up the
ramp?
Answer in units of m/s
2
.
1
Expert's answer
2020-01-06T10:11:32-0500

"m=42.4kg"

"a=9.81 m\/s2"

Considering the forces perpendicular to the inclined floor 


"mg*cos(10.1)=R+173*sin(26.7)"

solving above equation

"R=331.766 N"


FromF′=μR Where F' is the frictional force

"F\u2032=66.353 N (1)"


Applying "F=ma" along the inclined floor


"173*cos(26.7)-F'-mg*sin(10.1)=ma"

From (1)

"173*cos(26.7)-66.353-42.4*9.81*sin(10.1)=42.4a"


Solving above equation

"a=0.3598 m\/s2"